OpenStudy (mathhelp346):

PLEASE HELP PLEASE: Write each quadratic function in vertex form. Give the coordinates of the vertex and the equation of the axis of symmetry. Then describe the transformations from f(x)= x^2 to g. g(x)=-x^2+2

OpenStudy (anonymous):

do you know how to graph these two equation?

OpenStudy (mathhelp346):

no

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

can you see the vertex of each eq?

OpenStudy (mathhelp346):

..how do i enter them in?

OpenStudy (mathhelp346):

nevermind

OpenStudy (mathhelp346):

yes

OpenStudy (anonymous):

now that you know the vertex of each you can write the vertex form of each of the equation

OpenStudy (anonymous):

the vertex form formula is a(x-h)+k=0, where vertex= (h,k)

OpenStudy (mathhelp346):

oh ok

OpenStudy (mathhelp346):

so h is x and k is y

OpenStudy (anonymous):

from ax^2 +bx +c=0 f(x)=-x^2+2= x^2 where b=0, c=0 g(x)=-x^2+2, where b=0,c=2

OpenStudy (mathhelp346):

wait why is f(x)=-x^2+2= x^2? shouldn't it be just f(x)=x^2?

OpenStudy (anonymous):

these two eQ are different, see the graph

OpenStudy (mathhelp346):

why is the f(x) different?

OpenStudy (anonymous):

the main form is f(x)=x^2 but if you change them to y=-x^2 now theres a negative sign then if you add 2 then you have to jump up 2 on y axis

OpenStudy (anonymous):

did you see the difference between the two graph now?

OpenStudy (mathhelp346):

oh so your showing the transformation?